What mass of Br2 is needed to react with 9.7 g of aluminum? Round to two decimal places?



1
Expert's answer
2021-03-05T03:31:39-0500

2 Al + Br2 "\to" 2 Al Br

no. Of moles of Al = 9.7/26.98 = 0.359 moles

2 moles of Al react with 1 mole of Br2

0.359 moles of Al react with 0.1795 moles of Br2

Grams of Br2 = 0.1795×159.8 = 28.68g